OpenStudy (anonymous):

Graph the logarithmic function p(x) = -ln(x + 1) - 2. Find the asymptote. Then describe how the graph is transformed from the graph of its parent function. (How to graph a log and find asymptote?)

OpenStudy (anonymous):

answer: graphing calculator. Done. You're welcome!

OpenStudy (anonymous):

A neat little thing to know is that an asymptote occurs wherever the function cannot be defined. Since \(\ln(0)\) is not defined, then the asymptote must occur at \(x=-1\) since you'll have that \(p(-1)=-\ln(0)-2\).

OpenStudy (anonymous):

But a graphing calculator indeed helps with sketching the whole thing if you're not familiar with the variations of \(\ln(x)\). Just so you know, the parent function is the simplest general form of a function. So in this case, the parent function is \(\ln(x)\). For more, read here: http://en.wikipedia.org/wiki/Parent_function. Good luck!